OpenStudy (anonymous):

Evaluate the limit as x approaches -4. ((1/4)+(1/x))/(4+x) is the answer -1/16?

OpenStudy (anonymous):

@Chlorophyll is this right?

OpenStudy (cwrw238):

this simplifies to lim x--> -4 of 1/4x so i think thats right -1 /16 = -1/4 *(lim 1/x) = -1/16

OpenStudy (cwrw238):

i'm improving...

OpenStudy (anonymous):

yay thanks!

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):